OpenStudy (anonymous):

How to factor: x^2+2x+3x+6 Steps please :) Thanks!

OpenStudy (anonymous):

\[x^2+2x+3x+6\]

OpenStudy (mertsj):

(x+3)(x+2)

OpenStudy (anonymous):

(x-2)(x+4)

OpenStudy (anonymous):

Can you show me how you guys got that?

OpenStudy (anonymous):

yep

OpenStudy (mertsj):

\[x^2+2x+3x+6=x^2+5x+6=(x+3)(x+2)\]

OpenStudy (anonymous):

yep. what she said

OpenStudy (mertsj):

Or you could do it this way:

OpenStudy (mertsj):

\[x(x+2)+3(x+2)=(x+2)(x+3)\]

OpenStudy (mertsj):

That would be called factor by grouping.

OpenStudy (anonymous):

yaaa!